6. Reach is the number of different people exposed to a single medium, or in the case of a multimedia campaign, the entire media schedule. It may be expressed as the number of prospects or as a percentage of the target audience. For example if the target audience is 500,000 18- to 24-year-old males and 200,000 are exposed to the advertising, it may be expressed as a reach of 40 percent. Frequency is the number of times that each person in the audience is exposed to the media schedule. In this example, if the 200,000 men reached in the campaign generated 1 million exposures, the frequency would be 5.0: 1,000,000 ÷ 200,000 = 5.0 In buying media for our new toy, we are planning a 60 percent reach among kids A6-11. The U.S. Census report shows the population of that target at 22 million. Part 1: With that reach, how many kids would we then expose to the advertising? Insert Answer Here: Part 2: What is our frequency if the number of kids reached generated 39.6 million exposures?

(I do not need answer for number 6)

7. Reach and frequency multiplied together equal “gross rating points” or GRPs; for example, a 4-week reach of 50 at a frequency of 4 equals 200 GRPs. In our toy example (Question 6), what would our kids’ GRPs be? Insert Answer Here: a. 20 b. 200 c. 18 d. 180
Answer #1

Answer:

Part- 1:

Target Audience = 22 million

Planned Reach = 60%

Then, Expected No. of Kids Exposed to Advertising = 22 million * 60% = 13.2 million

Part-2:

Audience Exposed to Advertising = 13.2 million

Exposure Generated by Reach = 39.6 million

Then, Frequency = 39.6 / 13.2 = 3 times

Part-3;

GPRS = Percent of Target Market Reached * Exposure Frequency

Now, Reach of Advertising = 60% and Exposure Frequency = 3 times

Then, GPRS = 60 * 3 = 180